# Theodore S. Rappaport, Mingjun Ying, Nicola Piovesan, Antonio De Domenico, and Dipankar Shakya, “Waste Factor and Waste Figure: A Unified

### v1

@@ Abstract
This paper introduces Waste Factor (W), also denoted as Waste Figure (WF) in dB, a promising new metric for quantifying energy efficiency in a wide range of circuits and systems applications, including data centers and Radio Access Networks (RANs). Creating and managing 5G and future 6G networks that are energy-efficient is of paramount importance as the wireless industry evolves to become a major consumer of energy. Also, the networks used to connect data centers and artificial intelligence (AI) computing engines with users for machine learning (ML) applications must become more power efficient. This paper illustrates the limitations of existing energy efficiency metrics that inadequately capture the intricate energy dynamics of RAN components. We show here that W provides a generalized analysis of power utilization and power waste at both the component and system levels for any source-to-sink communication system. We delineate the methodology for applying W across various network configurations, including multiple-input single-output (MISO), single-input and multiple-output (SIMO), and multiple-input multiple-output (MIMO) systems, and demonstrate the effectiveness of W in identifying energy optimization opportunities. Our findings reveal that W not only offers nuanced insights into the energy performance of RANs but also facilitates informed decision-making for network design and operational efficiency. Furthermore, we show how W can be integrated with other key performance indicators (KPIs) to guide the development of optimal strategies for enhancing network energy efficiency under different operational conditions. Additionally, we present simulation results for a distributed multi-user MIMO (MU-MIMO) system at 3.5, 17, and 28 GHz, demonstrating overall network power efficiency on a per square kilometer basis, and show how overall W decreases (e.g., energy efficiency increases over the entire network) with an increasing number of base stations and increasing carrier frequency. This paper shows that adopting W as a figure of merit (FoM) can significantly contribute to the sustainability and energy optimization of next-generation wireless communication networks, paving the way for greener and more sustainable, energy-efficient 5G and 6G technologies.
@@
Waste Factor, Waste Figure, MIMO, RAN, energy efficiency, green communication, power consumption, sustainability, trade-off analysis.
@@ I INTRODUCTION
This paper introduces Waste Factor ($W$), or Waste Figure (WF) in dB, as a novel, powerful, and extensible metric to quantify energy efficiency within a cascade and for paralleled systems such as multiple-chain signal processing and antenna systems within RANs. Unlike traditional metrics, W provides a detailed perspective on how energy is utilized, wasted, or converted into useful output, thereby offering a comprehensive tool for energy analysis that can be applied far beyond just simple circuits or communication networks [1, 2, 3, 4]. The waste factor appears capable of quantifying the energy efficiency of any source-to-sink communication path, including computers, data centers, or other systems over which communications are carried out [2]. The Waste Factor theory is the first approach that we are aware of that breaks down the power consumption of any device or cascade into different components via superposition in order to characterize the work done for source-to-sink transport while considering the additive wasted power of such “on-path” components, and considering separately the work done by other components that are not even on the signal path of information flow. When such a dichotomy is made, it becomes an easy task to quantify energy usage and efficiency using the invariable fact that energy is neither created nor destroyed. That which is not delivered is wasted, and by such quantification as is done in the Waste Factor theory, it becomes possible to measure, model, and improve energy efficiency. Said another way, this theory lays bare the wasted and transported power and allows its modeling and optimization in a systematic way for any linear, cascaded, source-to-sink system.
In this paper, we examine energy efficiency within RANs in a comprehensive manner, emphasizing how W may be used as a means to compare and contrast energy efficiency within RAN systems and sub-systems. Our analysis compares the use of W with conventional energy efficiency metrics used in the standards bodies of today’s mobile industry, and demonstrates how Waste Factor provides valuable and nuanced insights into the energy dynamics of RAN components that are more useful than what is used today. We use W to evaluate the energy performance of various RAN elements and highlight its utility in pinpointing areas where energy optimization can yield significant benefits.
@@
Information and communication technology (ICT) infrastructures, such as data centers, mobile communication networks, and consumer devices, are essential in driving digital advancements and supporting various applications from the Internet to the Internet of Things (IoT) to the Internet and access to artificial intelligence (AI) engines. However, the ICT industry is a great consumer of energy, which incurs significant environmental costs. As of 2020, ICT infrastructures were responsible for consuming between 5-8% of global electricity production [5, 6] and were estimated to contribute between 1.5-4% of global greenhouse gas emissions [7, 8] and this was well before the ramp of AI. According to data from the International Telecommunication Union (ITU), two-thirds of the global population is now online [9]. Despite ongoing improvements in energy efficiency and the adoption of renewable energy sources, the environmental impact of the ICT sector is certain to increase to more than 20% of the global power consumption, driven by rising data traffic [1].
The introduction of 5G and the anticipation of 6G, both of which use bandwidths and deliver end-user data rates far beyond anything in history [10, 11], underline the urgency to address energy efficiency since it is well-known that greater consumption bandwidths necessitate larger energy consumption. The densification of networks, increased bandwidths, higher frequency bands, and enhanced data rates all exacerbate power consumption concerns. When considering the energy consumption of a typical cellular communications network, the RAN consumes 76% of the energy; the core network and data centers consume 19%, and other operations account for 5% of the total energy consumption (see Fig. 1[12, 13].
Many researchers have attempted to quantify and unify the analysis of energy consumption in mobile communications. In [14], the authors overview various energy efficiency metrics used in wireless systems at the component, equipment, and system/network levels. They note that energy efficiency metrics at the component and equipment levels are well-developed, while those for system/network levels need more attention. Authors in [14] note existing studies predominantly focus on macro-level energy efficiency metrics, often overlooking the intricate energy dynamics within RAN components. Such approaches do not fully capture the granular energy consumption patterns essential for devising targeted energy-saving strategies [15, 16]. To achieve long-term energy efficiency goals for the sustainable growth of the wireless industry, a unified metric is necessary to fill the gaps in previous energy efficiency metrics. In [17], McCune highlights that energy efficiency in wireless links varies by more than eleven orders of magnitude and stresses the need for a uniform approach to energy efficiency across different systems. In [18], the consumption factor (precursor to Waste Factor) evaluates the energy efficiency of a relay network and cascaded circuits. In [3], the consumption efficiency factor (CEF) provides a quantitative metric for the trade-off between the data rate and the power consumed by a communication system using Waste Factor and provides insights for network energy efficiency with different cell sizes. In the European Telecommunications Standards Institute (ETSI), Next Generation Mobile Networks (NGMN), ITU, and 3rd Generation Partnership Project (3GPP) standard bodies, energy efficiency (EE) metrics have been defined [19, 20, 12, 21, 22, 23].
@@
This paper considers the application of W for quantifying energy efficiency for different network configurations, including Multiple-Input Single-Output (MISO), Single-Input Multiple-Output (SIMO), and Multiple-Input Multiple-Output (MIMO) systems, and shows the versatility of W in assessing energy efficiency across a spectrum of network architectures. Additionally, we investigate the role of W in evaluating power efficiency within wired and wireless channels, offering a holistic approach to energy analysis in communication networks.
@@
Dynamic network conditions, such as fluctuating traffic demands, varying channel conditions, and diverse network topologies also benefit from a W-centered analysis to guide targeted enhancements, such as deploying energy-efficient hardware or implementing adaptive resource allocation algorithms, as well as long-term planning decisions, like network expansion or technology upgrade, to maintain high energy efficiency as the network evolves.
@@
The remainder of the paper is organized as follows. Section II defines W from theoretical underpinnings and highlights its advantages compared to conventional power efficiency metrics. Section III extends the Waste Factor theory to paralleled systems, such as MISO and MIMO. Section IV applies W to a 5G RAN case study and explains how W is used to determine power efficiency for both passive and active components. Section V emphasizes the utility of W in providing a comprehensive and nuanced perspective on energy efficiency in RANs, and illustrates the potential of W to inform decisions that can lead to more energy-efficient network operations than is currently found when using ETSI, NGMN, ITU or 3GPP energy efficiency metrics. Finally, Section VI concludes the paper by highlighting how the Waste Factor provides detailed insights into power utilization and energy efficiency optimization across various multiple-user wireless communication systems at different carrier frequencies, thus supporting the development of sustainable wireless networks.
@@ II-A2 Friis’s Formula for Noise factor
Here, $F_{i}$ is the noise factor and $G_{i}$ is the power gain of the $i^{th}$ device in the cascade where the first device is closest to the source, and the $N^{th}$ device is closest to the sink. Eq. (1) is essential for understanding how each component affects the overall noise performance of the system. It is readily apparent that the noise factor is primarily influenced by the input components with high gain, indicating that the initial component in the cascade, originating from the source side, exerts the most significant impact on the noise factor. The additional noise power contributed by a device in the cascade (independent of the input noise power) is quantified by
$P_{\text{additive-noise}}=(F-1)G_{i}N_{i}.$ (2)
@@ II-A3 Real-world Applications of Noise Factor
The practical importance of minimizing $F$ is highlighted in systems where maximum SNR is critical for extended link budget in low-power scenarios or for enabling extremely sophisticated multi-level digital modulations that are ultra-sensitive to SNR. $F$ facilitates informed decision-making regarding component selection and system design to achieve minimal noise impact and maximum performance. For example, in deep-space communication, even small increases in noise can lead to significant data loss [27]. In proposed 5G and 6G systems, Orthogonal frequency-division multiplexing (OFDM) with 4096-level quadrature amplitude modulation (QAM), which encodes 12 bits per symbol, requires SNR levels to exceed 22 dB when a coding rate of 1/3 is used under extended pedestrian A (EPA) fading mode to achieve a $10^{-2}$ bit error rate (BER) [28].
@@
With the move to wide bandwidth wireless channels (going from 20 MHz in 4G to 200 MHz in 5G), there is a need to allocate mobile services at higher carrier frequencies where spectrum is more plentiful [31, 32, 10, 11]. At such higher frequencies, the use of directional antennas on both ends of the link makes the channels more noise-limited and less interference-limited (the early cellular generations below 6 GHz were interference-limited [33]). As carrier frequencies and the commensurate channel bandwidths increase, the channels become more noise-limited or power-limited [10, 11, 33]. This shift brings about much greater power consumption due to the raw increase in consumed bandwidth and underscores the importance of optimizing power efficiency while also simultaneously minimizing noise for system sustainability. As shown here, W may become an essential metric in next-generation wireless communications for evaluating the power wasted in a cascade, which as shown here, is useful well beyond just receiver analysis, but may be used for understanding the power efficiency of any device, entire transmitters and receivers chains, and even the cascade between transmitters, receivers, and the propagation path loss of a channel.
@@ II-B Waste Factor
$W$ characterizes the power efficiency of a cascaded system by considering the power wasted by components along a cascade. Akin to $F$, the power wasted by a device/cascade can be modeled by observing the additive power waste, based on the output power at each successive stage, as the signal propagates from source to sink along the cascade. Such formulation provides an intuitive way to understand power waste at each cascade stage, as $W$ becomes a metric by which to compare the power efficiency of different devices/systems through their wasted power. A key concept in the development of $W$ is that wasted power is defined as the amount of power consumed by a component that carries the signal that is not found in the signal power output sent to the subsequent component in the cascade. In other words, the consumed energy that is not transmitted along the signal path cascade is classified as wasted energy. To derive $W$, the power consumption of an entire communication system between a source and a sink is represented by $P_{\text{consumed,total}}$. Then, it is noted that $P_{\text{consumed,total}}$ is equal to the superposition of the power $P_{\text{consumed,path}}$, which denotes the power consumed by all components along the signal path cascade, and for which $W$ is a useful parameter for computing, in addition to all of the power consumed by all components that are not on the path of the source-to-sink (this off-path power is denoted as $P_{\text{non-path}}$). To illustrate how W is used to compute the signal-path waste, which is only a portion of the total power consumed in any system, we break down the total consumed power into finer details. We observe first that $P_{\text{consumed,total}}$ may be split into four principal power components by superposition [18, 3, 30, 2]:
  • Input power from the source ($P_{\text{source, out}}$): the power input of the cascaded system that comes from the signal source or input side.
  • Output power ($P_{\text{signal}}$): the signal power delivered to the cascade output.
  • Non-signal path power ($P_{\text{non-signal}}$): the power consumed by devices within the cascade (e.g., along the signal path) to facilitate the signal transfer, but which is not part of the delivered signal power, and is viewed as wasted power which includes power consumed by active mixers, amplifiers, or power lost from heat dissipation and other losses in components along the signal path.
  • Non-path power ($P_{\text{non-path}}$) refers to the power consumed by components in the system that do not contribute to the signal power  [^1] Oscillators in electronic circuits are designed primarily for frequency synthesis, generating low-power periodic waveforms that serve as reference or control signals. They do not directly amplify or contribute power to the signal path, so their power consumption would be accounted for as non-path power. For amplifiers that are on the signal path, when they are turned off or not in use, the quiescent or standby power is shown in Section IV to be properly categorized as non-path power [2].. This includes the quiescent power of amplifiers that are not carrying signal, and the power used by oscillators, power dissipation units, displays, and other similar components not directly involved in the source-to-sink signal path.
@@
$=WP_{\text{signal}}+P_{\text{non-path}}.$ (4)
and where $P_{\text{system,added}}=P_{\text{signal}}-P_{\text{source,out}}$ is the power added (e.g., contributed) to the signal output power solely by the device or cascade.
@@
Note (3), (II-B), and (5) offer new ways to analyze, design, and compare the power efficiency of circuits and cascaded systems, since $W=1$ denotes no wasted power exists (all consumed power on the signal path is delivered as output signal power,) and $W=\infty$ denotes all power is wasted and no power is delivered to the output. Just like $F$, $W$ is generally defined regardless of the presence or level of signal input power applied by a signal source, and is only a function of the gain and the Waste factor of the device or system  [^2] As discussed in Section IV-IV-F, certain devices that transform between the analog and digital domains may use power levels to define $W$ as in (3)-(5), and we show in Section IV-IV-J that measurements of actual power consumption and power delivered may always be used to define W for a device or system using (3)-(5)..

### v2

# Waste Factor and Waste Figure: A Unified Theory for Modeling and Analyzing Wasted Power in Radio Access Networks for Improved Sustainability

@@ Abstract
This paper introduces Waste Factor (W), also denoted as Waste Figure (WF) in dB, a promising new metric for quantifying energy efficiency in a wide range of circuits and systems applications, including data centers and Radio Access Networks (RANs). Creating and managing 5G and future 6G networks that are energy-efficient is of paramount importance as the wireless industry evolves to become a major consumer of energy. Also, the networks used to connect data centers and artificial intelligence (AI) computing engines with users for machine learning (ML) applications must become more power efficient. This paper illustrates the limitations of existing energy efficiency metrics that inadequately capture the intricate energy dynamics of RAN components. We show here that W provides a generalized analysis of power utilization and energy waste at both the component and system levels for any source-to-sink communication system. We delineate the methodology for applying W across various network configurations, including multiple-input single-output (MISO), single-input and multiple-output (SIMO), and multiple-input multiple-output (MIMO) systems, and demonstrate the effectiveness of W in identifying energy optimization opportunities. Our findings reveal that W not only offers nuanced insights into the energy consumption of RANs but also facilitates informed decision-making for network design and operational efficiency. Furthermore, we show how W can be integrated with other key performance indicators (KPIs) and key value indicators (KVIs) to guide the development of optimal strategies for enhancing network energy efficiency under different operational conditions. Additionally, we present simulation results for a distributed multi-user MIMO (MU-MIMO) system at 3.5, 17, and 28 GHz, demonstrating overall network power efficiency on a per square kilometer basis, and show how overall W decreases (e.g., energy efficiency increases over the entire network) with an increasing number of base stations and increasing carrier frequency. This paper shows that adopting W as a figure of merit (FoM) can enable the design of more sustainable next-generation wireless communication networks, paving the way for greener and more sustainable, energy-efficient 5G and 6G technologies.
@@
Waste Factor, Waste Figure, MIMO, RAN, energy efficiency, green communication, power consumption, sustainability, KPI, KVI, trade-off analysis.
@@ I INTRODUCTION
Information and communication technology (ICT) infrastructures, such as data centers, mobile communication networks, and consumer devices, are essential in driving digital advancements and supporting various applications such as Internet of Things (IoT) and Artificial Intelligence (AI). However, the ICT industry is a great consumer of energy, which incurs significant environmental costs. As of 2020, ICT infrastructures were responsible for consuming between 5-8% of global electricity production [1, 2] and were estimated to contribute between 1.5-4% of global greenhouse gas emissions [3, 4] and this was well before the ramp of AI.
With two-thirds of the global population now online [5], and despite advances in energy efficiency and renewable energy adoption, the environmental impact of the ICT sector is poised to rise, potentially consuming over 20% of global power due to increased data traffic [6].
The introduction of 5G and the anticipation of 6G, both of which use bandwidths and deliver end-user data rates far beyond anything in history [7, 8], underline the urgency to address energy efficiency. Although 5G networks are about four times more energy-efficient per unit of data than 4G networks, their total energy consumption is considerably higher [9]. This increase is primarily due to the need for additional power-intensive processing to support wider bandwidths, more radio chains per base station, and increased cell site density to support higher frequency bands like C-band, mid-band, and millimeter wave [7, 10, 11, 12].
In the typical cellular network, the radio access network (RAN) is the largest power consumer, accounting for 76% of the total energy usage. The core network and data centers consume 19%, while other operations account for the remaining 5% of energy use, as depicted in Figure 1 [13, 14]. The financial implications of increased energy consumption are significant, particularly for mobile network operators. Network costs, which on average constitute about 25% of the operational expenditure (OPEX) of a mobile network operator, are heavily impacted by energy expenses, with approximately 90% of network costs dedicated to energy bills[15]. A nationwide carrier with 100 million subscribers pays about \$2 billion USD/year for electricity, and when multiplied over many hundreds of network operators (also called carriers) and 4 billion subscribers across the globe, the electricity bill for the cellular industry is approximately \$1 trillion USD/year. The challenges associated with reducing network energy consumption are exacerbated by a projected threefold increase in mobile data traffic from 2023 to 2029 [16]. Addressing this demand requires the deployment of additional infrastructure and innovative strategies to optimize energy usage.
@@
In response to the rising energy demands, the 3rd Generation Partnership Project (3GPP) has initiated efforts to mitigate RAN energy consumption. The 3GPP new radio (NR) Release 15 has introduced intra-NR network energy-saving solutions, including autonomous cell switch-off and re-activation features for capacity-boosting cells through Xn/X2 interfaces. Furthermore, 3GPP NR Release 17 has specified inter-system network energy-saving solutions and recognized network energy saving as a potential application for AI.
The trajectory of the network energy consumption growth underscores the urgent need for sustainable network expansion strategies and accurate, standardized energy efficiency metrics.
Many researchers have attempted to quantify and unify the analysis of energy consumption in mobile communications. In [17], the authors overview various energy efficiency metrics used in wireless systems at the component, equipment, and system/network levels. They note that energy efficiency metrics at the component and equipment levels are well-developed, while those for system/network levels need more attention. Authors in [17] note existing studies predominantly focus on macro-level energy efficiency metrics, often overlooking the intricate energy dynamics within RAN components. Such approaches do not fully capture the granular energy consumption patterns essential for devising targeted energy-saving strategies [18, 19]. To achieve long-term energy efficiency goals for the sustainable growth of the wireless industry, a unified metric is necessary to fill the gaps in previous energy efficiency metrics. In [20], McCune highlights that energy efficiency in wireless links varies by more than eleven orders of magnitude and stresses the need for a uniform approach to energy efficiency across different systems. In [21], the consumption factor evaluates the power consumption (e.g. the energy efficiency) of a relay network and cascaded circuits. In [22], the consumption efficiency factor (CEF) provides a quantitative metric for the trade-off between the data rate and the energy and power consumed by a communication system and provides insights for network energy efficiency with different cell sizes. In the European Telecommunications Standards Institute (ETSI), Next Generation Mobile Networks (NGMN), International Telecommunication Union (ITU), and 3GPP standard bodies, diverse energy efficiency (EE) metrics have been defined [23, 24, 13, 25, 26, 27].
This paper introduces Waste Factor ($W$), or Waste Figure (WF) in dB, as a novel, powerful, and extensible metric to quantify energy efficiency within a cascade and for paralleled systems such as multiple-chain signal processing and antenna systems within RANs. Unlike traditional metrics, W provides a detailed perspective on how energy is utilized, wasted, or converted into useful output, thereby offering a comprehensive tool for energy analysis that can be applied far beyond just simple circuits or communication networks [6, 28, 22, 29]. The waste factor is capable of quantifying the energy efficiency of any source-to-sink communication path, including computers, data centers, or other systems over which communications are carried out [28]. The Waste Factor theory is the first approach that we are aware of that breaks down the power consumption of any device or cascade into different components via superposition in order to characterize the work done for source-to-sink transport while considering the additive wasted power or squandered energy of such “on-path” components, and considering separately the work done by other components that are not even on the signal path of information flow. When such a dichotomy is made, it becomes an easy task to quantify energy usage and efficiency using the invariable fact that energy is neither created nor destroyed. That which is not delivered is wasted, and by using such quantification, as is done in the Waste Factor theory, it becomes possible to measure, model, and improve energy efficiency. Said another way, this theory lays bare the wasted and transported power and allows its modeling and optimization in a systematic way for any linear, cascaded, source-to-sink system.
This paper comprehensively examines energy efficiency within RANs, emphasizing how W can be used to compare and contrast energy efficiency within RAN systems and sub-systems. Our analysis compares the use of W with conventional energy efficiency metrics used in the standards bodies of today’s mobile industry such as ITU and ETSI, and demonstrates how W provides valuable and nuanced insights into the energy dynamics of RAN components. We use W to evaluate the energy performance of various RAN elements and highlight its utility in pinpointing areas where energy optimization can yield significant benefits.
@@
The paper also explores the application of W for quantifying energy efficiency in different network configurations, including Multiple-Input Single-Output (MISO), Single-Input Multiple-Output (SIMO), and Multiple-Input Multiple-Output (MIMO) systems, showcasing the versatility of W in assessing energy efficiency across a spectrum of network architectures.
@@
Dynamic network conditions, such as fluctuating traffic demands, varying channel conditions, and diverse network topologies, benefit from a W-centered analysis to guide targeted enhancements, such as deploying energy-efficient hardware or implementing adaptive resource allocation algorithms, as well as long-term planning decisions, like network expansion or technology upgrade, to maintain high energy efficiency as the network evolves.
@@
The remainder of the paper is organized as follows. Section II defines W from theoretical underpinnings and highlights its advantages compared to conventional energy efficiency and power efficiency metrics. Section III extends the Waste Factor theory to paralleled systems, such as MISO and MIMO. Section IV applies W to a 5G RAN case study and explains how W is used to determine power efficiency for both passive and active components. Section V emphasizes the utility of W in providing a comprehensive and nuanced perspective on energy efficiency in RANs, and illustrates the potential of W to inform decisions that can lead to more energy-efficient network operations than is currently found when using ETSI, NGMN, ITU or 3GPP energy efficiency metrics. Finally, Section VI concludes the paper by highlighting how the Waste Factor provides detailed insights into power utilization and energy efficiency optimization. We provide a realistic simulation comparing carrier frequency, antenna gain, and bandwidth considerations to demonstrate the practical use of W for determining power consumption for an operator, thereby illustrating how Waste Factor may be used to develop more sustainable wireless networks.
@@ II-A2 Friis’s Formula for Noise factor
Here, $F_{i}$ is the noise factor and $G_{i}$ is the power gain of the $i^{th}$ device in the cascade where the first device is closest to the source, and the $N^{th}$ device is closest to the sink. Eq. (1) is essential for understanding how each component affects the overall noise performance of the system. It is readily apparent from (1) that the noise factor of a cascade is primarily influenced by the input components with high gain, indicating that the initial component in the cascade, originating from the source side, exerts the most significant impact on the noise factor. The additional noise power contributed by the $i^{th}$ device in the cascade (independent of the input noise power) is defined by
$P_{\text{additive-noise,i}}=(F_{i}-1)G_{i}N_{i},$ (2)
@@ II-A3 Real-world Applications of Noise Factor
The practical importance of minimizing $F$ is critical for any communication system since improved SNR leads to better communication performance (e.g., lower bit error rate). Maximizing SNR is especially critical for extending the link budget in low-power scenarios or for enabling extremely sophisticated multi-level digital modulations that are ultra-sensitive to SNR. $F$ facilitates informed decision-making regarding component selection and system design to achieve minimal noise impact and maximum detector performance. For example, in deep-space communication, even small increases in noise can lead to significant data loss [33]. In proposed 5G and 6G systems, Orthogonal frequency-division multiplexing (OFDM) with 4096-level quadrature amplitude modulation (QAM), which encodes 12 bits per symbol, requires SNR levels to exceed 22 dB when a coding rate of 1/3 is used under the extended pedestrian A (EPA) fading mode to achieve a $10^{-2}$ bit error rate (BER) [34].
@@
With the move to wide bandwidth wireless channels (going from 20 MHz in 4G to 200 MHz or more in 5G), there is a need to allocate mobile services at higher carrier frequencies where spectrum is more plentiful [12, 10, 7, 8]. At such higher frequencies, the use of directional antennas on both ends of the link (e.g. dual directional channels) makes the channels more noise-limited and less interference-limited (the early cellular generations below 6 GHz were interference-limited [37]). As carrier frequencies and the commensurate channel bandwidths increase, the channels become more noise-limited or power-limited [7, 8, 37]. This paradigm shift to wider RF channel bandwidths brings about much greater power consumption due to the raw increase in consumed bandwidth and underscores the importance of optimizing power efficiency while also simultaneously minimizing noise for system sustainability. As shown here, W may become an essential metric in next-generation wireless communications for evaluating the power wasted in a cascade, which as shown here, is useful well beyond just receiver analysis, but may be used for understanding the power efficiency of any device, entire transmitter and receiver chains, and even the entire cascade between transmitters, receivers, and the propagation path loss of one or multiple channels.
@@ II-B Waste Factor
[…]
  • Input power from the source ($P_{\text{source, out}}$): the power applied to the device or the cascaded system that comes from the signal source or input side.
  • Output power ($P_{\text{signal}}$): the signal power delivered to the device or cascade output.
  • Non-signal path power ($P_{\text{non-signal}}$): the power consumed by devices within the cascade (e.g., along the signal path) to facilitate the signal transfer, but which is not part of the delivered signal power, and is viewed as wasted power which includes wasted power consumed by active mixers, amplifiers, or power lost from heat dissipation and other losses in components along the signal path.
  • Non-path power ($P_{\text{non-path}}$) refers to the power consumed by components in the system that do not contribute to the signal power. [^1] Oscillators in electronic circuits are designed primarily for frequency synthesis, generating low-power periodic waveforms that serve as reference or control signals, and are not on the direct signal path of a cascade, so their power consumption would be accounted for as non-path power. For amplifiers that are on the signal path, when they are turned off or not in use, the quiescent or standby power is shown in Section IV to be properly categorized as non-path power even though amplifiers are along the signal path during normal operation [28]. This includes the quiescent power of amplifiers that are not carrying signal, and the power used by oscillators, power dissipation units, displays, and other similar components not directly involved in the source-to-sink signal path.
@@
$=WP_{\text{signal}}+P_{\text{non-path}},$ (4)
where $P_{\text{system,added}}=P_{\text{signal}}-P_{\text{source,out}}$ is the power added (e.g., contributed) to the signal output power solely by the device or cascade with the applied signal source power removed from consideration under the assumption of linearity.
@@
Note (3), (II-B), and (5) offer new ways to analyze, design, and compare the power efficiency and energy efficiency of circuits and cascaded systems, since $W=1$ denotes no wasted power exists (all consumed power on the signal path is delivered as output signal power,) and $W=\infty$ denotes all power is wasted and no power is delivered to the output. Just like $F$, $W$ is generally defined regardless of the presence or level of signal input power applied by a signal source may be measured in-site, and is only a function of the gain and the Waste Factor of the device or system. [^2] As discussed in Section IV-IV-E, certain devices that transform between the analog and digital domains may use power levels to define $W$ as in (3)-(5), and we show in Section IV-IV-I that measurements of actual power consumption and power delivered may always be used to define W for a device or system using (3)-(5).